OpenStudy (anonymous):

What is the period of the function 2cos(3πx)?

OpenStudy (anonymous):

period in this case is equal to: Period= \[2pi/ 3pi = 2/3\]

OpenStudy (anonymous):

\[\frac{2\pi}{b}\] for period of \(\sin(bx)\) or \(\cos(bx)\)

OpenStudy (anonymous):

gd job satellite

OpenStudy (anonymous):

The factor 2 makes the curve bigger but doesn't affect the period, so just look at cos(3pi x). What is this value when x=0? When will the cosine again be that value?

OpenStudy (kropot72):

The angular frequency is 2*pi*f 2*pi*f = 3*pi*x from which f = (3/2)*x Period is 1/f = 2/(3*x)

OpenStudy (anonymous):

hey kropot72 what is the x at the bottom for?

OpenStudy (kropot72):

Because when the reciprocal of the expression was taken that's where it goes. If you want it at the top: Period T = (2*x^-1)/3

OpenStudy (anonymous):

that is not what I mean, look at this and check where it talks about the period, you do not need the x at all. http://science.kennesaw.edu/~plaval/applets/TrigCos.html

OpenStudy (kropot72):

At your link the period is a function of the variable k. Note that k is the denominator. In the question here the variable has been named x.